A loan of $10,000 is to be repaid in 10 years according to the one of the following arrangements:

(a) monthly installments of $A in arrears with an APR of 8%,

(b) monthly payments of $B in arrears with an APR of 9.1%, with a cashback of $684 at the beginning, i.e. the loan amount is reduced by that amount,

(c) monthly payments of $C in arrears with a quoted interest rate of 7.2% on annual rest.

Calculate the sum A+B+C
Answer #1

317 318 10000 Loan Amount Period(in year) 319 320 (a) APR 321 322 323 (b) 324 Monthly Installment APR cashback Monthly Instal

Formulas Used :-

Monthly installment (a) =PMT(J321,J319,-J318)/12

Monthly installment (b)=PMT(J323,J319,-J318+J324)/12

Monthly installment (c)=PMT(J326,J319,-J318)/12

short summary :-

in our calculations we have used Excel and particularly used the PMT function of Excel which provide us the value of monthly installment and but we have used the data annually and at last we have devided whole function by 12 so we get monthly installment for the case (b) where Cashback has given we have subtract the amount of cash back from Amount of total loan and all other calculation remain the same.

I hope my efforts will be fruitful to you....?